# Hopf Bifurcation of Nonlinear Non-Hermitian Skin Effect

### v1

@@ Abstract
Topological phases of matter constitute a cornerstone in condensed matter physics. Specifically, in band insulators and Bogoliubov-de Gennes superconductors, topology of wave functions has been extensively classified [1, 2, 3]. The nontrivial bulk topology manifests itself in the emergence of anomalous gapless states at boundaries: bulk-boundary correspondence. Beyond quantum materials, topological band theory extends its applicability to a broad spectrum of classical synthetic platforms, including photonic [4, 5] and mechanical [6, 7] systems, as well as active matter [8]. Importantly, in such classical settings, nonlinearity is an inherent and pervasive feature. Accordingly, the impact of nonlinearity on topological insulators has attracted growing interest in recent years [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. For example, nonlinearity has been shown to induce solitons in topological insulators [9, 10, 11, 13, 14]. It has also been demonstrated to quantize a topological pump through soliton formation and pitchfork or saddle-node bifurcations [20, 22, 23].
@@
Meanwhile, topological characterization of non-Hermitian systems has also been studied considerably in both theory [31, 32, *Esaki-11, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] and experiments [49, 50, *Zhou-18, 52, 53]. Non-Hermiticity generally arises from the exchange of particles and energy with the external environment and gives rise to a variety of phenomena that have no counterparts in conservative systems [54, 55]. The bulk-boundary correspondence for non-Hermitian topology manifests itself as the skin effect, anomalous localization of a macroscopic number of eigenstates driven by nonreciprocal dissipation [56, 57, 58, 59, *YSW-18-Chern, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. Given that such an extraordinary sensitivity to boundary conditions is absent in Hermitian systems, the skin effect is a universal phenomenon intrinsic to non-Hermitian systems. It has been observed across various open classical experiments of mechanical systems [80, *Ghatak-19-skin-exp], electric circuits [82, *Hofmann-19-skin-exp, 84], photonic systems [85], and active matter [86], in addition to open quantum ones of single photons [87], ultracold atoms [88, 89], and digital processors [90].
@@
Notably, both nonlinearity and non-Hermiticity comprise important ingredients underlying the physics of open systems. Prime recent examples include topological lasers, where the high-efficiency lasing is accomplished through the interplay of non-Hermiticity, topology, and nonlinearity [91, 92, 93, 94, 95, 96, *Bandres-18]. Consequently, the fate of topological phases in the simultaneous presence of nonlinearity and non-Hermiticity has recently been investigated [98, 99, 100, *Yuce-25, 102, 103, 104, 105, *ManyManda24soliton, 107, 108]. Nonlinear non-Hermitian eigenvalue problems are also relevant to the Green’s function formalism at equilibrium [109] and boost deformations [110, *Guo-23]. However, the role of nonlinearity on the non-Hermitian skin effect, along with associated unique nonlinear non-Hermitian phenomena, has remained largely unexplored.
@@
In this Letter, we reveal the Hopf bifurcation of the nonlinear skin effect as a critical phenomenon unique to nonlinear non-Hermitian systems. Introducing a nonlinear extension of the Hatano-Nelson model, a prototypical model exhibiting the skin effect in the linear regime [112, *Hatano-Nelson-97], we demonstrate that nonlinearity destabilizes skin states and instead gives rise to the emergence of delocalized states. We further show that this localization transition of the nonlinear skin effect corresponds to the Hopf bifurcation in the theory of dynamical systems [114, 115]—a transition from a stable fixed point to a limit cycle in phase space. Moreover, we uncover that the exponential localization of skin states transforms into the algebraic (i.e., power-law) localization precisely at the Hopf bifurcation point, serving as a hallmark of the criticality. We also substantiate these behavior in a lattice counterpart of the nonlinear Hatano-Nelson model. Our work elucidates a significant role of nonlinearity in the skin effect and provides distinctive dissipative phenomena arising from the interplay of non-Hermiticity and nonlinearity.
@@
where $E$ and $\psi$ denote an eigenvalue and its eigenstate, respectively. Specifically, we study a nonlinear extension of the Hatano-Nelson model [112, *Hatano-Nelson-97],
@@
with a momentum operator $\hat{k}\coloneqq-\text{i}\partial_{x}$. Here, $\gamma\in\mathbb{R}$ ($\varepsilon\geq 0$) quantifies the degree of linear (nonlinear) non-Hermiticity. The non-Hermitian term $\gamma>0$ enhances right-propagating waves with positive momenta and suppresses left-propagating ones with negative momenta, manifesting as the asymmetry of the hopping dynamics in real space. While this asymmetric hopping physically describes a tilted magnetic field in type-II superconductors for the original Hatano-Nelson model [112, *Hatano-Nelson-97], it can also be implemented, for example, in active mechanical [80, *Ghatak-19-skin-exp], electric [82, *Hofmann-19-skin-exp, 84], and photonic [85] systems. Meanwhile, $\varepsilon$ characterizes the modulation of the asymmetric hopping that depends nonlinearly on the local wave function amplitude $\left|\psi\right|^{2}$. We assume that the norm of wave functions is unconstrained, corresponding to open classical systems.
@@
with $k=\pi n/L$ ($n\in\mathbb{N}_{+}$). Notably, the spectra exhibit substantial differences depending on the boundary conditions, which is a hallmark of the skin effect unique to non-Hermitian systems. Accordingly, all the wave functions are localized toward the right (left) for $\gamma>0$ ($\gamma<0$).
Figure 2: Hopf bifurcation in the nonlinear Hatano-Nelson model ($\varepsilon=0.3$, $E=0.5$) for (a, c) $\gamma=-0.2$ and (b, d) $\gamma=0.2$. (a, b) Flow in phase space $\left(\psi,\partial_{x}\psi\right)$. (c, d) Wave function numerically obtained with the initial conditions $\psi\left(x=0\right)=0$, $\psi^{\prime}\left(x=0\right)=1$.
Hopf bifurcation.—We now proceed to study the influence of nonlinearity $\varepsilon>0$. As discussed earlier, the linear non-Hermitian term with $\gamma<0$ induces an exponential decay of the wave function toward the right, making the nonlinear term $\varepsilon\left|\psi\right|^{2}$ negligible. For $\gamma>0$, by contrast, the linear non-Hermitian term amplifies the wave function whereas the nonlinear term gives a negative feedback mechanism. Their competition can give rise to the formation of a delocalized state unique to nonlinear non-Hermitian systems. We demonstrate that this is indeed the case and that nonlinearity destabilizes skin states and leads to delocalized states corresponding to a limit cycle in phase space. Analogous to the linear regime, we below focus on the case of $E>0$ and $\psi\in\mathbb{R}$. Notably, by replacing space with time, the nonlinear Schrödinger equation (1) governing the nonlinear Hatano-Nelson model in Eq. (2) reduces to the equation of motion for the van der Pol oscillator [116]—a prototype exhibiting a limit cycle.
@@
and consider the associated flow in phase space $\left(\psi,\partial_{x}\psi\right)$. We then study the spatial distribution of the wave function $\psi$ toward the right direction $x\geq 0$, corresponding to the semi-infinite boundary conditions. In the linear regime $\varepsilon=0$, the system exhibits a unique fixed point at the origin $\left(\psi,\partial_{x}\psi\right)=\left(0,0\right)$. Around this fixed point, the eigenvalues of the coefficient matrix are
@@
Even in the nonlinear regime, the origin $\left(\psi,\partial_{x}\psi\right)=\left(0,0\right)$ remains to be the unique fixed point, which is characterized by the eigenvalues in Eq. (6) and thus stable (unstable) for $\gamma<0$ ($\gamma>0$) [Fig. 2 (a, b)]. A crucial distinction from the linear case is the emergence of a limit cycle for $\gamma>0$: a flow departing around the origin does not diverge to infinity but instead asymptotically approaches a closed periodic orbit. Consequently, while the wave function decays toward the right for $\gamma<0$ similar to the linear regime [Fig. 2 (c)], it undergoes spatial delocalization for $\gamma>0$, despite the presence of the asymmetric hopping [Fig. 2 (d)]. This transition from a stable fixed point to a limit cycle represents the Hopf bifurcation in the framework of dynamical systems. We here unveil that the Hopf bifurcation manifests itself as the localization transition of the nonlinear skin effect.
Limit cycle.—For $\gamma>0$, the existence and uniqueness of a limit cycle in phase space $\left(\psi,\partial_{x}\psi\right)$ can be proved based on the Liénard theorem [117, 118, 119, 120]. Accordingly, delocalized eigenstates generally appear for $\gamma>0$ even in the presence of the asymmetric hopping. Since their emergence is caused by the relative competition between $\gamma$ and $\varepsilon$, strong nonlinearity is not necessarily needed, which enables a perturbative approach. Indeed, in the weakly nonlinear regime [$0<\varepsilon\ll 1$, $\gamma=\mathcal{O}\left(\varepsilon\right)$], the energy dispersion of these delocalized eigenstates is obtained as
@@
on the basis of singular perturbation theory. The distinction from the linear case in Eq. (4) manifests itself in the additional constant term. Conversely, in the strongly nonlinear regime ($\varepsilon\gg 1$), the energy dispersion is evaluated as
@@
Notably, despite the nonrelativistic nature of the original system, the strong nonlinearity induces the emergent relativistic energy dispersion.
@@
Even in the nonlinear regime $\varepsilon\neq 0$, plane waves $\psi\left(x\right)=ae^{\text{i}kx}$ remain to be eigenstates with the eigenenergies $E\left(k\right)=k^{2}/2+\text{i}\,(\gamma-\varepsilon\left|a\right|^{2})\,k$ under the periodic boundary conditions. As a result, a point gap persists except for $\left|a\right|^{2}\neq\gamma/\varepsilon$, in which case the winding number is determined as $W\left(E_{\rm P}\right)=\mathrm{sgn}\,(\gamma-\varepsilon\left|a\right|^{2})$ for $E_{\rm P}>0$. This holds even for $\gamma>0$, where nonlinearity destabilizes skin states and leads to the emergence of delocalized states characterized by limit cycles. Thus, the correspondence between the point-gap topology and skin effect seems to break down in the nonlinear regime. Nevertheless, the emergence of delocalized eigenstates associated with limit cycles may still be related to point-gap topology in nonlinear non-Hermitian systems. It should also be noted that nonlinearity can give rise to additional eigenstates beyond plane waves in contrast to the linear regime. Such additional eigenstates may restore the bulk-boundary correspondence for point-gap topology, which is left for future studies.
Figure 3: Nonlinear Hatano-Nelson model on a lattice ($E=0.5$). Initial conditions: $\psi_{0}=0$, $\psi_{1}=1$. (a-c) Wave function for $\varepsilon=0.3$ and (a) $\gamma=-0.2$, (b) $\gamma=0.2$, and (c) $\gamma=0$ (dotted line: $\left|\psi_{n}\right|=3/\sqrt{n}$). (d) Inverse of the localization length, $1/\xi$ ($\left|\psi_{n}\right|\sim e^{n/\xi}$), numerically obtained over $500\leq n\leq 1000$ [$\varepsilon=0$ (gray), $\varepsilon=0.3$ (red)].
Lattice model.—We further elucidate the Hopf bifurcation of the nonlinear non-Hermitian skin effect in a lattice counterpart of Eq. (2). A discrete counterpart of the nonlinear Schrödinger equation on a one-dimensional lattice is given by $\hat{H}\left(\ket{\psi}\right)\ket{\psi}=E\ket{\psi}$ with
$|\displaystyle\hat{H}\left(\ket{\psi}\right)=\sum_{n}\left(-\frac{1-\gamma+\varepsilon\left|\psi_{n}\right|^{2}}{2}\ket{n+1}\bra{n}\right.$ |
$\qquad\left.-\frac{1+\gamma-\varepsilon\left|\psi_{n}\right|^{2}}{2}\ket{n}\bra{n+1}+\ket{n}\bra{n}\right),$ (12)
where |$\ket{\psi}=\left(\psi_{1},\cdots,\psi_{L}\right)^{T}$ represents a single-particle eigenstate, and |$\ket{n}$ denotes a single-particle state localized at site $n$. Similar to the continuum counterpart, $\gamma\in\mathbb{R}$ characterizes the degree of asymmetric hopping, and $\varepsilon\geq 0$ quantifies the strength of its nonlinear modulation that can induce a negative feedback mechanism for $\gamma>0$. Such a tight-binding model can be implemented, for example, in a laser-written array of waveguides, where individual elements are coupled with each other evanescently [30].
To investigate the localization behavior, we here employ the transfer matrix method [121]. We reformulate the nonlinear Schrödinger equation with the nonlinear transfer matrix $T\left(\psi_{n}\right)$:
@@
The spatial distributions of wave functions |$\ket{\psi}$ for given energy $E$ are thus determined through the cumulative product of $T\left(\psi_{n}\right)$’s:
$\begin{pmatrix}\psi_{L+1}\\ \psi_{L}\end{pmatrix}=\left[\prod_{n=1}^{L}T\left(\psi_{n}\right)\right]\begin{pmatrix}\psi_{1}\\ \psi_{0}\end{pmatrix}.$ (15)
@@
Specifically, the localization length is characterized by $1/\xi=\lim_{L\to\infty}\alpha_{L}/L$, where $\alpha_{L}$ denotes the smallest singular value of $\prod_{n=1}^{L}T\left(\psi_{n}\right)$. A similar nonlinear transfer matrix approach has recently been used in the Hermitian setting [27, 28].
@@
We numerically calculate wave functions and their localization properties (Fig. 3). Consistent with the continuum counterpart, the wave functions are localized toward the left for $\gamma<0$ but become delocalized for $\gamma>0$. Precisely at the Hopf bifurcation point $\gamma=0$, the wave function exhibits the algebraic decay $\left|\psi_{n}\right|\propto 1/\sqrt{n}$, in agreement with Eq. (10). Figure 3 (d) shows the localization length $\xi$ as a function of linear non-Hermiticity $\gamma$. In the linear regime (i.e., $\varepsilon=0$), we obtain $\xi=1/\gamma$, indicating the conventional skin effect. In the nonlinear regime, on the other hand, while the skin effect persists for $\gamma<0$, the localization length diverges, $\xi=\infty$, for $\gamma>0$, implying the emergence of delocalized states. These results demonstrate the Hopf bifurcation and accompanying delocalization of skin states even in the lattice model.
@@
Discussion.—In this Letter, we have elucidated the Hopf bifurcation of the nonlinear skin effect as a phase transition induced by the interplay of non-Hermiticity and nonlinearity. Since this phase transition is governed by the relative competition between linear and nonlinear non-Hermitian terms, strong nonlinearity is not strictly required, thereby potentially facilitating its experimental implementation. While we have focused on a nonlinear extension of the Hatano-Nelson model in this Letter, the Hopf bifurcation is a ubiquitous phenomenon in nonlinear non-Hermitian systems, driving phase transitions in the nonlinear skin effect. Further exploration of the implications of other types of bifurcation for the nonlinear skin effect is warranted.
@@
Moreover, the effect of disorder merits further investigation. We find the emergence of nonlinearity-induced mobility edges even in the absence of disorder [120], reminiscent of those found in the disordered Hatano-Nelson model [112, *Hatano-Nelson-97]. The combination of disorder and nonlinearity can form distinctive mobility edges. Developing a unified understanding of the localization transitions in nonlinear non-Hermitian systems can also bear practical significance for disorder-resilient wave transport phenomena. Additionally, the skin effect can be enriched in different symmetry classes or higher dimensions in the linear regime [70, 72, 73, 75]. Understanding the influence of nonlinearity therein is worth studying. Finally, the nonlinearity-induced instability of the skin effect implies that nonlinearity can fundamentally change the topological classification and bulk-boundary correspondence.

### v2

@@ Abstract
Topological phases of matter constitute a cornerstone in condensed matter physics. In band insulators and Bogoliubov-de Gennes superconductors, topology of wave functions has been extensively classified [1, 2, 3]. The nontrivial bulk topology manifests itself in the emergence of anomalous gapless states at boundaries: bulk-boundary correspondence. Beyond quantum materials, topological band theory extends its applicability to a broad spectrum of classical synthetic platforms, including photonic [4, 5] and mechanical [6, 7] systems, as well as active matter [8]. Importantly, in such classical settings, nonlinearity is an inherent and pervasive feature. Accordingly, the impact of nonlinearity on topological insulators has attracted growing interest in recent years [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. For example, nonlinearity has been shown to induce solitons in topological insulators [9, 10, 11, 13, 14]. It has also been demonstrated to quantize a topological pump through soliton formation and pitchfork or saddle-node bifurcations [20, 22, 23].
@@
Meanwhile, topological characterization of non-Hermitian systems has also been studied considerably in both theory [31, 32, *Esaki-11, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] and experiments [49, 50, *Zhou-18, 52, 53]. Non-Hermiticity generally arises from the exchange of particles and energy with the external environment and gives rise to a variety of phenomena that have no counterparts in conservative systems [54, 55]. The bulk-boundary correspondence for non-Hermitian topology appears as the skin effect, anomalous localization of a macroscopic number of eigenstates driven by nonreciprocal dissipation [56, 57, 58, 59, *YSW-18-Chern, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79]. Given that such an extraordinary sensitivity to boundary conditions is absent in Hermitian systems, the skin effect is a universal phenomenon intrinsic to non-Hermitian systems. It has been observed across various open classical experiments of mechanical systems [80, *Ghatak-19-skin-exp], electric circuits [82, *Hofmann-19-skin-exp, 84], photonic systems [85], and active matter [86], in addition to open quantum ones of single photons [87], ultracold atoms [88, 89], and digital processors [90].
@@
Notably, both nonlinearity and non-Hermiticity comprise important ingredients underlying the physics of open systems. Prime recent examples include topological lasers, where the high-efficiency lasing is accomplished through the interplay of non-Hermiticity, topology, and nonlinearity [91, 92, 93, 94, 95, 96, *Bandres-18]. Consequently, the fate of topological phases in the simultaneous presence of nonlinearity and non-Hermiticity has recently been investigated [98, 99, 100, *Yuce-25, 102, 103, 104, 105, 106, *ManyManda24soliton, 108, 109, 110, 111]. Nonlinear non-Hermitian eigenvalue problems are also relevant to the Green’s function formalism at equilibrium [112] and boost deformations [113, *Guo-23]. The combination of the skin effect and nonlinearity was shown to induce the single-mode lasing [103] and soliton formation [107, 110]. However, the role of nonlinearity on the non-Hermitian skin effect, along with associated unique nonlinear non-Hermitian phenomena, has remained largely unexplored.
@@
In this Letter, we reveal the Hopf bifurcation of the nonlinear skin effect as a critical phenomenon unique to nonlinear non-Hermitian systems. Introducing a nonlinear extension of the Hatano-Nelson model, a prototypical model exhibiting the skin effect in the linear regime [115, *Hatano-Nelson-97], we demonstrate that nonlinearity destabilizes skin states and instead gives rise to the emergence of delocalized states. We further show that this localization transition of the nonlinear skin effect corresponds to the Hopf bifurcation in the theory of dynamical systems [117, 118]—a transition from a stable fixed point to a limit cycle in phase space. Moreover, we uncover that the exponential localization of skin states transforms into the algebraic (i.e., power-law) localization precisely at the Hopf bifurcation point, serving as a hallmark of the criticality. We also substantiate these behavior in a lattice counterpart of the nonlinear Hatano-Nelson model. Our work elucidates a significant role of nonlinearity in the skin effect and provides distinctive dissipative phenomena arising from the interplay of non-Hermiticity and nonlinearity.
@@
where $E$ and $\psi$ denote an eigenvalue and its eigenstate, respectively. Specifically, we study a nonlinear extension of the Hatano-Nelson model [115, *Hatano-Nelson-97],
@@
with a momentum operator $\hat{k}\coloneqq-\text{i}\partial_{x}$. Here, $\gamma\in\mathbb{R}$ ($\varepsilon\geq 0$) quantifies the degree of linear (nonlinear) non-Hermiticity. The non-Hermitian term $\gamma>0$ enhances right-propagating waves with positive momenta and suppresses left-propagating ones with negative momenta, manifesting as the asymmetry of the hopping dynamics in real space. While this asymmetric hopping physically describes a tilted magnetic field in type-II superconductors for the original Hatano-Nelson model [115, *Hatano-Nelson-97], it can also be implemented, for example, in active mechanical [80, *Ghatak-19-skin-exp], electric [82, *Hofmann-19-skin-exp, 84], and photonic [85] systems. Meanwhile, $\varepsilon$ characterizes the modulation of the asymmetric hopping that depends nonlinearly on the local wave function amplitude $\left|\psi\right|^{2}$. Under the nonreciprocal operation $\hat{k}\mapsto-\hat{k}$, this model returns back to itself only if both $\gamma$ and $\varepsilon$ simultaneously change their signs. We assume that the norm of wave functions is unconstrained, corresponding to open classical systems.
@@
with $k=\pi n/L$ ($n\in\mathbb{N}_{>0}$). Notably, the spectra exhibit substantial differences depending on the boundary conditions, which is a hallmark of the skin effect unique to non-Hermitian systems. Accordingly, all the wave functions are localized toward the right (left) for $\gamma>0$ ($\gamma<0$).
Figure 2: Hopf bifurcation in the nonlinear Hatano-Nelson model ($\varepsilon=0.3$, $E=0.5$) for (a, c) $\gamma=-0.2$ and (b, d) $\gamma=0.2$. (a, b) Flow in phase space $\left(\psi,\partial_{x}\psi\right)$. The gray loop shows the limit cycle. (c, d) Wave function numerically obtained with the boundary conditions $\psi\left(x=0\right)=0$, $\psi^{\prime}\left(x=0\right)=1$.
Hopf bifurcation.—We now proceed to study the influence of nonlinearity $\varepsilon>0$. As discussed earlier, the linear non-Hermitian term with $\gamma<0$ induces an exponential decay of the wave function toward the right, making the nonlinear term $\varepsilon\left|\psi\right|^{2}$ negligible. For $\gamma>0$, by contrast, the linear non-Hermitian term amplifies the wave function whereas the nonlinear term gives a negative feedback mechanism. Their competition can give rise to the formation of a delocalized state unique to nonlinear non-Hermitian systems. We demonstrate that this is indeed the case and that nonlinearity destabilizes skin states and leads to delocalized states corresponding to a limit cycle in phase space. Below, we focus on the case of $E>0$ and $\psi\in\mathbb{R}$ to elucidate the influence of nonlinearity on linear skin states. Notably, by replacing space with time, the nonlinear Schrödinger equation (1) governing the nonlinear Hatano-Nelson model in Eq. (2) reduces to the equation of motion for the van der Pol oscillator [119]—a prototype exhibiting a limit cycle.
@@
and consider the associated flow in phase space $\left(\psi,\partial_{x}\psi\right)$. We then investigate the spatial distribution of the wave function $\psi$ toward the right direction $x\geq 0$, corresponding to the semi-infinite boundary conditions, and study how the localization behavior depends on the eigenenergy $E$. In the linear regime $\varepsilon=0$, the system exhibits a unique fixed point at the origin $\left(\psi,\partial_{x}\psi\right)=\left(0,0\right)$. Around this fixed point, the eigenvalues of the coefficient matrix are
@@
Even in the nonlinear regime, the origin $\left(\psi,\partial_{x}\psi\right)=\left(0,0\right)$ continues to serve as the unique fixed point, which is characterized by the eigenvalues in Eq. (6) and thus stable (unstable) for $\gamma<0$ ($\gamma>0$) [Fig. 2 (a, b)]. A crucial distinction from the linear case is the emergence of a limit cycle for $\gamma>0$: a flow departing around the origin does not diverge to infinity but instead asymptotically approaches a closed periodic orbit. Consequently, while the wave function decays toward the right for $\gamma<0$ similar to the linear regime [Fig. 2 (c)], it undergoes spatial delocalization for $\gamma>0$ [Fig. 2 (d)]. This transition from a stable fixed point to a limit cycle represents the Hopf bifurcation in the framework of dynamical systems. We here unveil that the Hopf bifurcation manifests itself as the localization transition of the nonlinear skin effect.
Limit cycle.—For arbitrary $\gamma>0$, the existence and uniqueness of a limit cycle in phase space $\left(\psi,\partial_{x}\psi\right)$ are proved based on the Liénard theorem [120, 121, 122, 123]. Accordingly, delocalized eigenstates generally appear for $\gamma>0$ even in the presence of the asymmetric hopping. Hence, their emergence does not necessitate strong nonlinearity, which enables a perturbative approach. Indeed, in the weakly nonlinear regime $0<\varepsilon\ll 1$, $\gamma=\mathcal{O}\left(\varepsilon\right)$], the energy dispersion of these delocalized eigenstates is obtained as
@@
on the basis of singular perturbation theory (see the End Matter and Supplemental Material [123] for details). Notably, all the eigenstates with $E\geq\gamma^{2}/4$ are delocalized, in contrast to the skin effect. The distinction from the linear case in Eq. (4) is also reflected in the additional constant term. Conversely, in the strongly nonlinear regime ($\varepsilon\gg 1$), the energy dispersion is evaluated as
@@
Despite the nonrelativistic nature of the original system, the strong nonlinearity induces the emergent relativistic energy dispersion.
@@
Even in the nonlinear regime $\varepsilon\neq 0$, plane waves $\psi\left(x\right)=ae^{\text{i}kx}$ persist as eigenstates with the eigenenergies $E\left(k\right)=k^{2}/2+\text{i}\,(\gamma-\varepsilon\left|a\right|^{2})\,k$ under the periodic boundary conditions. As a result, a point gap persists except for $\left|a\right|^{2}\neq\gamma/\varepsilon$, in which case the winding number is determined as $W\left(E_{\rm P}\right)=\mathrm{sgn}\,(\gamma-\varepsilon\left|a\right|^{2})$ for $E_{\rm P}>0$. This holds even for $\gamma>0$, where nonlinearity destabilizes skin states and leads to the emergence of delocalized states characterized by limit cycles. Thus, the correspondence between the point-gap topology and skin effect seems to break down in the nonlinear regime. Nevertheless, the emergence of delocalized eigenstates associated with limit cycles may still be related to point-gap topology in nonlinear non-Hermitian systems. It should also be noted that nonlinearity can give rise to additional eigenstates beyond plane waves in contrast to the linear regime. Such additional eigenstates may restore the bulk-boundary correspondence for point-gap topology, which is left for future studies.
Figure 3: Nonlinear Hatano-Nelson model on a lattice ($E=0.5$). Boundary conditions: $\psi_{0}=0$, $\psi_{1}=1$. (a-c) Wave function for $\varepsilon=0.3$ and (a) $\gamma=-0.2$, (b) $\gamma=0.2$, and (c) $\gamma=0$ (dotted line: $\left|\psi_{n}\right|=3/\sqrt{n}$). (d) Inverse of the localization length, $1/\xi$ ($\left|\psi_{n}\right|\sim e^{n/\xi}$), numerically obtained over $500\leq n\leq 1000$ $\varepsilon=0$ (gray), $\varepsilon=0.3$ (red)].
[…]
To investigate the localization behavior under the semi-infinite boundary conditions, we here employ the transfer matrix method [124]. We reformulate the nonlinear Schrödinger equation with the nonlinear transfer matrix $T\left(\psi_{n}\right)$:
@@
[…]
$\begin{pmatrix}\psi_{L+1}\\ \psi_{L}\end{pmatrix}=\left[\prod_{n=1}^{L}T\left(\psi_{n}\right)\right]\begin{pmatrix}\psi_{1}\\ \psi_{0}\end{pmatrix}\eqqcolon T_{L}\begin{pmatrix}\psi_{1}\\ \psi_{0}\end{pmatrix}$. (15)
@@
Specifically, the localization length is characterized by $1/\xi=\lim_{L\to\infty}\alpha_{L}/L$, where $\alpha_{L}$ denotes the smallest positive eigenvalue of $\log\sqrt{T_{L}^{{\dagger}}T_{L}}$. A similar nonlinear transfer matrix approach has recently been used in the Hermitian setting [27, 28].
@@
We numerically calculate wave functions and their localization properties under the boundary conditions $\psi_{0}=0$ and $\psi_{1}=1$ (Fig. 3). Consistent with the continuum counterpart, the wave functions are localized toward the left for $\gamma<0$ but become delocalized for $\gamma>0$. Precisely at the Hopf bifurcation point $\gamma=0$, the wave function exhibits the algebraic decay $\left|\psi_{n}\right|\propto 1/\sqrt{n}$, in agreement with Eq. (10). Figure 3 (d) shows the localization length $\xi$ as a function of linear non-Hermiticity $\gamma$. In the linear regime (i.e., $\varepsilon=0$), we obtain $\xi=1/\gamma$, indicating the conventional skin effect. In the nonlinear regime, on the other hand, while the skin effect persists for $\gamma<0$, the localization length diverges, $\xi=\infty$, for $\gamma>0$, implying the emergence of delocalized states. These results demonstrate the Hopf bifurcation and accompanying delocalization of skin states even in the lattice model. The Hopf bifurcation also manifests in the nonlinear dynamics, providing a clear experimental signature: whereas a wave packet propagates unidirectionally in the linear regime, nonlinearity causes the bidirectional propagation even in the presence of the asymmetric hopping, reflecting the nonlinearity-induced destabilization of the skin effect and the concomitant emergence of delocalized states (see the End Matter for details).
@@
Discussion.—In this Letter, we have elucidated the Hopf bifurcation of the nonlinear skin effect as a phase transition induced by the interplay of non-Hermiticity and nonlinearity. Since this phase transition is governed by the relative competition between linear and nonlinear non-Hermitian terms, strong nonlinearity is not strictly required, thereby potentially facilitating its experimental implementation. While we have focused on a nonlinear extension of the Hatano-Nelson model in this Letter, the Hopf bifurcation is a ubiquitous phenomenon in nonlinear non-Hermitian systems, driving phase transitions in the nonlinear skin effect. Indeed, we further demonstrate it in a nonlinear non-Hermitian extension of the Su-Schrieffer-Heeger model [125, 123]. Further exploration of the implications of other types of bifurcation for the nonlinear skin effect is warranted.
@@
Moreover, the effect of disorder merits further investigation. We find the coexistence of delocalized and localized eigenstates in the lattice version of the nonlinear Hatano-Nelson model in Eq. (12[123]. This shows the emergence of nonlinearity-induced mobility edges even in the absence of disorder, reminiscent of those found in the disordered Hatano-Nelson model [115, *Hatano-Nelson-97]. The combination of disorder and nonlinearity can form distinctive mobility edges. Developing a unified understanding of the localization transitions in nonlinear non-Hermitian systems can also bear practical significance for disorder-resilient wave transport phenomena. Additionally, the skin effect can be enriched in different symmetry classes or higher dimensions in the linear regime [70, 72, 73, 75]. Understanding the influence of nonlinearity therein is worth studying. Finally, the nonlinearity-induced instability of the skin effect implies that nonlinearity can fundamentally change the topological classification and bulk-boundary correspondence.